Shift report 20000718 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jul 18 2000 - 16:06:42 EDT

  • Next message: http account: "Shift report 20000719 16:00-24:00"

    ------------------------------------------------------------
    Date: Tue Jul 18 16:06:41 2000, Shift: 08:00-16:00
    Supervisor: Michael Murray, with: Yuri 
    Log book#3 pages 26 - 27
    
    Shift Summary:
    ------------------------------------------------------------
    There was no useful beam on this shift.  RHIC was still 
    doing RAMP studies. Flemming updated SuperMon to handel
    the new Low gain ZDC ADC. There seems to be some 
    problem with the attenuators used. Michael is working
    on this. Dana replaced a "paddle card" for the TOFW. That
    improved things but still there are problems. Ramero
    started the flow of isobutane in T3 and T5.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Jul 18 2000 - 16:10:04 EDT